A recent college graduate is weighing two job offers in the sales industry. Company #1 will pay $35,000 per year with 5% commission while company #2 has offered $40,000 with 4% commission. How much would the salesman need to sell in order for him to make more money working for company #1?


1
Expert's answer
2021-05-07T11:45:42-0400

Given data


Company 1                                                                                      Company 2

Fixed pay = $35,000                                                                   fixed pay = $40,000

Variable pay = 5% commission                                               variable pay = 4% commission


Let ‘x’ be the amount of sales done by the person

The amount earned in company 1 = fixed pay + variable pay

                                                             = 35,000 + (5% of x)

                                                            = 35,000 + ( "\frac{5}{100}x)"

                                                            = 35,000 + 0.05x

The amount earned in company 2 = fixed pay + variable pay

                                                            = 40,000 + (4% of x)

                                                            = 40,000 + ("\frac{4}{100}x)"

                                                            = 40,000 + 0.04x

Given that the amount earned in company 1 should be greater than amount earned in company 2

From the above statement we can write the equation as

 35,000 + 0.05x ≥ 40,000 + 0.04x

0.05x – 0.04x ≥ 40,000 – 35,000

 0.01x ≥ 5,000

x ≥ 5,00,000


To earn more in company 1 than in company 2 he should sell total amount greater than $5,00,000